the state of chiapas mexico decided to fund a program for improving reading skills in elementary school students. the first cost is 325000 now, and an update of 170000 every 6 years forever. determine the perpetual equivalent annual cost at an interest rate of 10% per year
Answer #1

First cost = 325,000

Annual cost = 325000*10% = 32,500

170,000 every 6 years needs to be annualised to find the annual amount for a perpetuity.

Since the first outgo will be after 6 years,

Annual amount = 170,000*(F/A, 10%, 6) = 22,033.26

Perpetual equivalent annual cost = 32,500+22,033.26 = 54,033.26